# See or Say Graphs: Agent-Driven Scalable Graph Understanding with Vision-Language Models

### v1

@@ H.2 Complex Local Tasks: Visual Modality with Visual Graph Thoughts
Task 1: Node Existence and Properties Prompt Template: "Determine if node <node_id> exists in the graph. If so, what is its degree?" GraphVista Behavior: The planning agent categorizes this as a simple property task. The model performs a targeted retrieval query against the GraphRAG Base $\mathcal{K}$ to find the entry for <node_id>. The retrieved context, containing the node’s 1-hop neighborhood information, is then used to generate a direct answer. Rationale: This task involves retrieving explicitly stored attributes. The textual representation in the GraphRAG base is optimized for such direct lookups, making it efficient and reliable.
@@
Task 2: Edge Existence and Weight Prompt Template: "Is there a direct edge between node <node_id_1> and node <node_id_2>? If yes, provide its weight." GraphVista Behavior: The query is identified as a simple property task. The system retrieves the adjacency information for <node_id_1> from $\mathcal{K}$. The VLM then processes this structured text to check for a connection to <node_id_2> and extracts the corresponding weight attribute. Rationale: Like node property lookups, edge verification relies on retrieving specific, pre-processed information. The textual modality avoids the potential for visual ambiguity (e.g., overlapping edges or nodes) that could occur in a complex graph visualization.
@@
Task 3: Shortest Path Identification Prompt Template: "Find and list the sequence of nodes that form the shortest path between node <node_id_1> and node <node_id_2>." GraphVista Behavior: The planning agent categorizes this as a complex local task. The Visual Graph Thoughts Agent extracts a multi-centric subgraph centered around the shortest path candidates between <node_id_1> and <node_id_2>. This high-resolution subgraph is visualized. The VLM then initiates a Visual Graph Thought process, highlighting nodes and edges step-by-step in the image to trace the optimal path, verbalizing its reasoning at each step before presenting the final sequence. Rationale: Textual descriptions of paths can be convoluted and difficult for VLMs to reason over effectively. A visual representation provides a “what-you-see-is-what-you-get” view of the topology, which is more intuitive for complex structural reasoning and naturally supports stepwise inference grounded in visual evidence.

### v2

@@ I.2 Visual-Modality Tasks: Reasoning with Visual Graph Thoughts
Task 1: Node Existence and Properties Prompt Template: "Determine if node <node_id> exists in the graph. If so, what is its degree?" GraphVista Behavior: The planning agent categorizes this as a text-modality task. The model performs a targeted retrieval query against the GraphRAG Base $\mathcal{K}$ to find the entry for <node_id>. The retrieved context, containing the node’s 1-hop neighborhood information, is then used to generate a direct answer. Rationale: This task involves retrieving explicitly stored attributes. The textual representation in the GraphRAG base is optimized for such direct lookups, making it efficient and reliable.
@@
Task 2: Edge Existence and Weight Prompt Template: "Is there a direct edge between node <node_id_1> and node <node_id_2>? If yes, provide its weight." GraphVista Behavior: The query is identified as a text-modality task. The system retrieves the adjacency information for <node_id_1> from $\mathcal{K}$. The VLM then processes this structured text to check for a connection to <node_id_2> and extracts the corresponding weight attribute. Rationale: Like node property lookups, edge verification relies on retrieving specific, pre-processed information. The textual modality avoids the potential for visual ambiguity (e.g., overlapping edges or nodes) that could occur in a complex graph visualization.
@@
Task 3: Shortest Path Identification Prompt Template: "Find and list the sequence of nodes that form the shortest path between node <node_id_1> and node <node_id_2>." GraphVista Behavior: The planning agent categorizes this as a visual-modality task. The Visual Graph Thoughts Agent extracts a multi-centric subgraph centered around the shortest path candidates between <node_id_1> and <node_id_2>. This high-resolution subgraph is visualized. The VLM then initiates a Visual Graph Thought process, highlighting nodes and edges step-by-step in the image to trace the optimal path, verbalizing its reasoning at each step before presenting the final sequence. Rationale: Textual descriptions of paths can be convoluted and difficult for VLMs to reason over effectively. A visual representation provides a “what-you-see-is-what-you-get” view of the topology, which is more intuitive for complex structural reasoning and naturally supports stepwise inference grounded in visual evidence.
@@
Task 5: Graph Diameter Calculation (Modality-Collaborative) Prompt Template: "Calculate the diameter of the graph. First identify potential peripheral nodes, and then determine the maximum shortest path distance between them." GraphVista Behavior: The Planning Agent identifies this as a collaborative task and decomposes it into two sequential sub-tasks. Phase 1 (Text Modality): The agent queries the GraphRAG Base $\mathcal{K}$ to identify a set of "peripheral nodes" (typically defined as nodes with low centrality scores in Tier 3). This filters the search space to the most likely candidates for the diameter endpoints. Phase 2 (Visual Modality): The system extracts multi-centric subgraphs connecting these candidate pairs. The Visual Graph Thoughts agent then visually traces the shortest paths between them to calculate their distances. Finally, the system aggregates these results and returns the maximum distance found as the graph diameter. Rationale: Calculating diameter requires both global search and local pathfinding. The text modality is efficient for globally filtering candidate nodes based on statistical properties (centrality), avoiding the need to process the entire graph visually. Conversely, the visual modality excels at the subsequent topological reasoning required to accurately count path lengths between specific pairs, which is cumbersome for text-based reasoning.